OpenStudy (anonymous):

Here's a problem that is asking for me to simplify that I can not figure out: (3x^2/y)^-3 (-2y^-3/3y^4)^2

OpenStudy (amistre64):

the -3 exponent means to flip the contents

OpenStudy (amistre64):

or simply use it as is and remember the rules of multiplying exponsnts

OpenStudy (amistre64):

3^(-3) - 1/3^3 = 1/27 x^2.-3 = x^-6 = 1/x^6 1/y^-3 = y^3

OpenStudy (amistre64):

... = 1/3^3....typoes it lol